1) Check cable between RFPD and FSS box for quality. Replace with a good short cable.
2) Using a directional coupler, look at the RFPD output in lock on a scope with 50 Ohm term.
I suspect its a lot of harmonics because we're overmodulating to compensate for the bad
3) Purchase translation stages for the FSS mode matching lenses. Same model as the PMC lenses.
Fix the mode matching.
4) Get the shop to build us up some more bases for the RFPDs on the PSL such as we have for the LSC.
Right now they're on some cheesy Delrin pedestals. Too soft...
5) Dump the beam reflected off the FSS RFPD with a little piece of black glass or a razor dump.
Anodized aluminum is no good and wiggles too much.
Your PZT is broken.
you asked for: diff 2008/07/21,13:00 2008/07/22,2:44:16 utc '.*FSS.*|.*MC.*'
LIGO controls: differences, 2008 07/21 13:00:00 utc vs. 2008 07/22 02:44:16 utc
__Epics_Channel_Name______ __Description__________ __value1____ __value2____
C1:IOO-MC_REFL_GAIN 22.000000 19.000000
C1:IOO-MC_REFL_OFFSET 0.818380 0.818100
C1:PSL-FSS_MGAIN 10.000000 30.000000
C1:PSL-FSS_PHCON 2.073170 1.413170
Component Schematic Actual
--------- --------- ------
C140 10u open
C144 10u open
C149 open a gray Cap. value unknown
C141 10u open
C145 10u open
R97 1.58K 0
R99 open 1130
R103 open 1130
R100 open 0
R104 100 1130
R98 1.58K open
R109 367 365
I unlocked the PMC and swept over C1:PSL-PMC_RAMP's full range a couple of times this morning. The PMC should now be relocked and returned
BLUE: Output of filter bank during an FM9 turn off. This is the transient which goes to the DAC.
The transients are mostly of the same magnitude as this.
RED: This is the input of the filter module during another such transient.
GREEN: Tried another switch; this time I filtered the time series in DTT by typing the SimDW into
the Triggered Time Series filter field. This should be simulating what comes out of the
output of the DW board - to convert to volts multiply by 15/32768.
PURPLE: Same kind of filtering as the GREEN, but with also a double 30 Hz highpass to remove the low
frequency damping control signals. You can see that the total transient is only ~5 counts
or ~1 mV at the coil driver output. This is comparable to the relative offset in the bypass
and filter paths.